# Dense Backpropagation Improves Training for Sparse Mixture-of-Experts

### v2

@@ Abstract
Mixture of Experts (MoE) pretraining is more scalable than dense Transformer pretraining, because MoEs learn to route inputs to a sparse set of their feedforward parameters. However, this means that MoEs only receive a sparse backward update, leading to training instability and suboptimal performance. We present a lightweight approximation method that gives the MoE router a dense gradient update while continuing to sparsely activate its parameters. Our method, which we refer to as Default MoE, substitutes missing expert activations with default outputs consisting of an exponential moving average of expert outputs previously seen over the course of training. This allows the router to receive signals from every expert for each token, leading to significant improvements in training performance. Our Default MoE outperforms standard TopK routing in a variety of settings without requiring significant computational overhead. Code: https://github.com/vatsal0/default-moe.
Figure 1: Pretraining Benchmarks for Default MoE. Our Default MoE improves MoE pretraining by providing a default vector for unactivated expert activations. After training for 320B tokens on FineWeb-Edu (Lozhkov et al., 2024), our 2B parameter Default MoE outperforms a standard TopK MoE on various pretraining benchmarks. All evals are run with the lm-eval-harness (Gao et al., 2024).
@@ 1 Introduction
Large-scale pretraining hinges on scaling up the number of parameters in a model, because models with more parameters are more sample-efficient and require less training to reach the same performance as smaller models (Kaplan et al., 2020; Hoffmann et al., 2022). The most common model architecture is a dense Transformer architecture (Vaswani et al., 2023) because its performance scales well with parameters and data. However, a sparsely activated Mixture-of-Experts (MoE) Transformer architecture (Shazeer et al., 2017) has been used by many industry deployments (DeepSeek-AI Team, 2024b; xAI, 2024; Databricks, 2024; Mistral Team, 2024; Snowflake, 2024; DeepSeek-AI Team, 2024a) because MoEs have been shown to scale even better than dense Transformers (Clark et al., 2022; Du et al., 2022; Lepikhin et al., 2020; Fedus et al., 2022). MoEs learn a routing function that selectively activates the Top-K subset of their modules, or experts, most relevant to a given input. This conditionally sparse activation (Jacobs et al., 1991; Jordan & Jacobs, 1994) allows us to multiplicatively increase the model parameter count without significantly increasing the cost of training or inference.
At train time, the sparse router enables very large MoEs to be trained with relatively little computation per token, but it also presents a challenge. The router does not receive a gradient update from experts that it does not activate. This may slow down learning, as the gradient update is unable to adjust the router to promote the optimal expert for each token. It may also contribute to load imbalance – where a few experts are over-utilized – leading to inefficient training and resource usage (Zoph et al., 2022; Zhou et al., 2022). This can be corrected by activating all experts during training, enabling all experts and routing directions to be optimized at once, although this requires extreme computational costs.
In this work we propose a new method that strikes a middle ground between dense and sparse routing; our proposed router can receive and differentiate contributions from all experts, while the computational cost remains virtually identical to a standard Top-K router. In our proposed method, the router receives contributions from all experts for every token. However, a forward pass is only computed on the Top-K experts for each token, while other experts contribute a “default vector” that represents the expected output of a typical token. Our method adds minimal computational overhead compared to a standard Top-K router while significantly improving model training speed and performance.
@@ 2 Background & Related Work
The Router Gradient. Consider the gradient of the MoE layer’s output $y$ with respect to the router parameters $W$. We express $y$ as a function of $W$ by combining Equation 1 and Equation 2. With the chain rule, we can backpropagate through this function by considering the gradient at each respective step:
@@
The second term in Equation 3, $\frac{\partial\pi}{\partial W}$, is straightforward to compute because the steps in Equation 1 are easily differentiable, as they consist of linear operations and activations. But the first term, $\frac{\partial y}{\partial\pi}$, is not differentiable because in Equation 2 Top-K expert selection transforms the continuous router weights $\pi\in\mathbb{R}^{N}$ into a discrete set of selected experts $\mathcal{A}$ with $\binom{N}{K}$ possible values. One way to address backpropagation of nondifferentiable operations is to use the straight-through estimator (Bengio et al., 2013), which treats the operation as the identity function. With straight-through we bypass the Top-K routing function and Equation 2 becomes the dot product between $\pi$ and the vector of all $E_{i}(x)$ with the following gradient:
@@
Figure 2: Overview of Routing with Dense Approximations. The original mixture of experts router only receives gradients corresponding to experts the token is routed to, because there is no output from other experts. Our approach provides the router with a complete (dense) gradient by letting non-activated experts contribute a default vector that approximates its output without the cost of a forward pass. As indicated by the dashed green arrows, the approximated gradients are not actually connected to the token in the computation graph; instead, they are artificially applied in the backward pass.
Related Works. Previous work has tried to address the issue of routing in MoEs. Separate from Top-K is the Sinkhorn routing method (Clark et al., 2022).  Fedus et al. (2022) propose an auxiliary loss that encourages load balancing.  Dai et al. (2024) propose multiple additional auxiliary loss terms.  Wang et al. (2024) propose learning biases rather than an auxiliary load balancing loss, and this is used in DeepSeek-AI Team (2024b). Phi-3.5-MoE (Phi Team, 2024) uses SparseMixer (Liu et al., 2024, 2023), another estimator for ${\partial y}/{\partial\pi}$ not involving straight-through. We provide a comparison between our approach and SparseMixer later in Section 4.3. Recently, ReLU routing (ReMoE) has been proposed as an alternative to the traditional method of TopK followed by softmax (Wang et al., 2025). Similar to our method, ReMoE addresses the non-differentiable nature of the sparse MoE router. Our approach is to still use straight-through and auxiliary loss with a TopK router, but approximate these additional expert outputs. We now present our method.
@@ 3 Designing Dense Backpropagation
In this section we design a new MoE router that can receive a dense gradient without requiring additional forward passes through experts. Equation 4 describes the dense gradient of the router for an individual token. Each expert output $E_{i}(x)$ in the router gradient corresponds to updating the embedding for expert $i$ in the routing layer (i.e. the $i$th row of the router parameters $W$). In a standard sparse MoE, the gradient only includes expert outputs $E_{i}$ for the experts $i\in\mathcal{A}$ selected by the Top-K routing function. The other terms are essentially 00, which means that non-activated experts can not influence the router’s update for a token. Our goal is to approximate the dense gradient in Equation 4 without directly computing the missing terms $E_{i^{\prime}}(x)$ for $i^{\prime}\notin\mathcal{A}$. This enables the router to receive a signal from all experts in its gradient update, without actually activating all of these experts. The router can then consider information from all experts when learning to route tokens instead of being limited to only receiving feedback from experts that were activated. As a result, we expect an improved router which learns to allocate tokens to different experts more effectively, leading to improved MoE training performance.
Figure 2 presents an overview of our method for updating the router. Since we wish to avoid additional computational overhead, we fill in the missing terms in Equation 4 with estimates $\hat{E_{i}}$ for non-activated experts $E_{i}$. This simple addition ensures that the router updates its weights for all experts instead of only those selected by the Top-K routing.
@@ 3.1 Approximating Missing Expert Outputs
Backpropagating through the sparse MoE activation in Equation 2 does not align with the router’s true dense gradient specified in Equation 4. Specifically, the gradient in TopK routing leads to an error based on the non-activated experts. Let us consider the gradient of the router parameters $W$ with respect to the loss $\mathcal{L}$. For clarity, we examine the gradient term for a single input $x$; our approach also translates to practically training the router with a batch of inputs.
Using the chain rule, we decompose the router gradient to isolate the nondifferentiable $\frac{\partial y}{\partial\pi}$, similar to Equation 3:
@@
The error term corresponding to $\frac{\partial y}{\partial\pi}$, $\sum_{i\notin\mathcal{A}}(E_{i}(x)-\mathbb{E}_{x}[E_{i}(x)])$, is 00 in expectation for all $i$. As a result, our default vector method corrects the gradient error by filling in missing expert activations with the mean expert output.
@@ 3.2 Default Expert Outputs with Exponential Moving Averages
To enable dense backpropagation without sacrificing the computational efficiency of sparse MoE forward passes, we propose Default MoE. The Default MoE provides a default value for the outputs of non-selected experts using exponential moving averages (EMAs). The EMA accounts for the fact that experts are being updated, and thus their outputs for inputs $x$ will change over the course of training. For each expert $E_{i}$, we maintain an EMA of its average output:
@@
This approach enables dense backpropagation through the router while preserving the sparse computational pattern that makes MoE architectures efficient. The router receives gradient information about all possible routing decisions, not just the selected experts, potentially leading to more informed routing decisions.
@@ 4 Evaluation
We describe the experimental setup in Section 4.1, show that our Default MoE outperforms the TopK baseline in Section 4.2, and then show how our method compares to the baseline as we ablate virtually every parameter in the experimental setup in Section 4.3. We finish by concluding that our method does not significantly impact throughput or the memory footprint of training, so the improvements we observe are -as far as we can tell- a free lunch.
@@ 4.1 Experimental Setup
Model Architectures. We use NcK to refer to a model with $N$ total experts and $K$ active experts. For our main results in Figure 1Figure 4Figure 4 we train a standard MoE with 8 experts. All of our models have 1.96 billion total parameters, 366 million of which are non-MoE parameters. This leaves 1.6B MoE parameters, and the number of active parameters depends on sparsity. For example, a 32 expert MoE with 2 active experts has sparsity 1/16, so 100M of the MoE parameters are active i.e. the 32c2 MoE has 466M active parameters. All parameter counts are specified in Table 1.
@@
We ablate the model architecture, hidden dim, number of total experts, and number of active experts in Section 4.3.
Dataset. We train on FineWeb-Edu (Lozhkov et al., 2024) (for Figure 1 and Figure 4) and FineWeb (Penedo et al., 2024) (for all other plots and results) with the Llama3 tokenizer (Llama 3 Team, 2024). We use SwiGLU (Shazeer, 2020) MLPs following Llama (Touvron et al., 2023), $16$ attention heads with dimension of $64$, LayerNorm (Ba et al., 2016) and RoPE (Su et al., 2023).
Hyperparameters. We use the initialization from  Wang et al. (2022) for residual branch merge layers and the initialization from Nguyen & Salazar (2019) for all other layers. We use the AdamW optimizer (Loshchilov & Hutter, 2019). All hyperparameters for the model that produces Figure 1 can be found in our config file, and all hyperparameters are the same between the baseline and our DefaultMoE.
MoE-Specific Training Details. We set the auxiliary loss (Fedus et al., 2022) to $0.01$. We do not use z-loss or jitter, because we find that at this scale they do not improve training for the baseline model. Following DeepSeek-AI Team (2024b), we set the first layer to be dense. Following Liu et al. (2024), we compute the aux loss across nodes. We train dropless MoEs.
Implementation. We train with the gpt-neox library (Andonian et al., 2023) integrated with Megablocks (Gale et al., 2022) and augmented with Triton kernels from (Hsu et al., 2024). We open-source our training code.
@@ 4.2 Main Results
Model MathQA LogiQA MMLU OpenBookQA Lambada SocialIQA ARC HellaSwag Winogrande PIQA Average
Score Random Baseline 20 25 25 25 1.6111We use the “Random word from passage” as a baseline for LAMBADA (Paperno et al., 2016). 33.3 25 25 50 50 -
TopK MoE 25.8 26.1 31.8 37.2 38.8 39.5 45.9 52.3 53.5 71.1 42.2
Default MoE 25.9 27.2 32.5 38.4 41.3 40.9 47.8 53.3 54.7 71.7 43.4
Difference +0.4% +4.2% +2.2% +3.2% +6.4% +3.5% +4.1% +1.9% +2.2% +0.8% +2.8%
Improvement Over Random TopK MoE 5.8 1.1 6.8 12.2 37.2 6.2 20.9 27.3 3.5 21.1 14.2
Default MoE 5.9 2.2 7.5 13.4 39.7 7.6 22.8 28.3 4.7 21.7 15.3
Difference +1.7% +100% +10.3% +9.8% +6.7% +22.6% +9.1% +3.7% +30.6% +2.8% +7.5%
Table 2: Comparison of Pretraining Benchmark Scores for TopK MoE and Default MoE. After training for 320 billion tokens, Default MoE outperforms TopK MoE on all benchmarks, with an average improvement of 2.8%. We also report the difference normalized by the random baseline. For example, Default MoE scores 7.5% higher than the random baseline on MMLU compared to TopK’s 6.8%, for a relative improvement of 10.3%.
Pretraining Benchmarks. Our main result compares our Default MoE, which performs a dense update of the router weights by approximating the dense gradient, to a baseline Top-K MoE, both with $K=1$ active experts. After training both our Default MoE and a TopK MoE for 320 billion tokens on FineWeb-Edu (Lozhkov et al., 2024), we evaluate both models on a variety of common pretraining benchmarks in Table 2. We plot benchmarks for both models in Figure 1. Our Default MoE outperforms TopK MoE across the board, with an average benchmark improvement of 2.8%.
In Figure 4 we compare each MoE’s loss curves. Without introducing significant overhead, our method reduces the tokens required to reach a target perplexity of $12.18$ by $9\%$. As the model has just $500M$ active parameters, we are training for significantly more than the compute-optimal number of tokens Hoffmann et al. (2022). Therefore, we can be confident that we are not just seeing our method converge faster to a worse minima; this is a real improvement in terms of both speed of convergence and the quality of the converged model.
@@
In Figure 1 and Figure 4 we pretrain on FineWeb-Edu (Lozhkov et al., 2024). We also pretrain models on FineWeb (Penedo et al., 2024), which is less curated. The rest of the plots in this paper (our ablations) are on FineWeb, in order to enforce that on some level we are not simply overfitting to the pretraining dataset that we use to collect benchmark scores. In Figure 4 we show that when using the learning rate of $7\times 10^{-4}$ that we sweep in Figure 8, our DefaultMoE outperforms the baseline TopK. Therefore, our improvements are not due to any unfair hyperparameter tuning when compared to the baseline.
Default MoE Beats Sparsemixer.  Liu et al. (2023, 2024) use Sparsemixer, which estimates the true router gradient without straight-through.  Liu et al. (2024) note that Sparsemixer lags behind Top-K (which our method always outperforms) for the first $0.5T$ tokens, likely due to the noise that Sparsemixer adds. It is somewhat outside of our computational budget to run a baseline for $0.5T$ tokens. We see in Figure 5 that our method significantly outperforms SparseMixer for at least the first $10B$ tokens.

### v3

@@ Abstract
Mixture of Experts (MoE) pretraining is more scalable than dense Transformer pretraining, because MoEs learn to route inputs to a sparse set of their feedforward parameters. However, this means that MoEs only receive a sparse backward update, leading to training instability and suboptimal performance. We present a lightweight approximation method that gives the MoE router a dense gradient update while continuing to sparsely activate its parameters. Our method, which we refer to as Default MoE, substitutes missing expert activations with default outputs consisting of an exponential moving average of expert outputs previously seen over the course of training. This allows the router to receive signals from every expert for each token, leading to significant improvements in training performance. Our Default MoE outperforms standard TopK routing in a variety of settings without requiring significant computational overhead.
@@ 1 Introduction
The sparsely activated Mixture-of-Experts (MoE) Transformer architecture (Shazeer et al., 2017) has been used by many industry deployments (DeepSeek-AI Team, 2024b; xAI, 2024; Databricks, 2024; Mistral Team, 2024; Snowflake, 2024; DeepSeek-AI Team, 2024a) because MoEs have been shown to scale even better than dense Transformers (Clark et al., 2022; Du et al., 2022; Lepikhin et al., 2020; Fedus et al., 2022). MoEs learn a routing function that selectively activates the Top-K subset of their modules, or experts, most relevant to a given input. This conditionally sparse activation (Jacobs et al., 1991; Jordan and Jacobs, 1994) allows us to multiplicatively increase the model parameter count without significantly increasing the cost of training or inference. At train time, the sparse router enables very large MoEs to be trained with relatively little computation per token, but it also presents a challenge. The router does not receive a gradient update from experts that it does not activate. This may slow down learning, as the gradient update is unable to adjust the router to promote the optimal expert for each token. This can be corrected by activating all experts during training, enabling all experts and routing directions to be optimized at once, although this requires much more compute.
In this work we propose DefaultMoE: a new method that strikes a middle ground between dense and sparse routing; our proposed router can receive and differentiate contributions from all experts, while the computational cost remains virtually identical to a standard Top-K router. In our proposed method, the router receives contributions from all experts for every token. However, a forward pass is only computed on the Top-K experts for each token, while other experts contribute a “default vector" that represents the expected output of a typical token. Our method adds minimal computational overhead compared to a standard Top-K router while improving performance on a range of standard benchmarks for 2 billion total parameter MoEs trained on 160 billion tokens.
@@ 2 Background & Related Work
The Router Gradient. Consider the gradient of the MoE layer’s output $y$ with respect to the router parameters $W$. We express $y$ as a function of $W$ by combining Equation˜1 and Equation˜2. With the chain rule, we can backpropagate through this function by considering the gradient at each respective step:
@@
The second term in Equation˜3, $\frac{\partial\pi}{\partial W}$, is straightforward to compute because the steps in Equation˜1 are easily differentiable, as they consist of linear operations and activations. But the first term, $\frac{\partial y}{\partial\pi}$, is not differentiable because in Equation˜2 Top-K expert selection transforms the continuous router weights $\pi\in\mathbb{R}^{N}$ into a discrete set of selected experts $\mathcal{A}$ with $\binom{N}{K}$ possible values. One way to address backpropagation of nondifferentiable operations is to use the straight-through estimator (Bengio et al., 2013), which treats the operation as the identity function. With straight-through we bypass the Top-K routing function and Equation˜2 becomes the dot product between $\pi$ and the vector of all $E_{i}(x)$ with the following gradient:
@@
Figure 1: Overview of Routing with Dense Approximations. The original MoE router only receives gradients corresponding to experts the token is routed to, because there is no output from other experts. Our approach provides the router with a complete (dense) gradient by letting non-activated experts contribute a default vector that approximates its output without the cost of a forward pass. As indicated by the dashed green arrows, the approximated gradients are not actually connected to the token in the computation graph; instead, they are artificially applied in the backward pass.
@@ 3 Designing Dense Backpropagation
In this section we design a new MoE router that can receive a dense gradient without requiring additional forward passes through experts. Equation˜4 describes the dense gradient of the router for an individual token. Each expert output $E_{i}(x)$ in the router gradient corresponds to updating the embedding for expert $i$ in the routing layer (i.e. the $i$th row of the router parameters $W$). In a standard sparse MoE, the gradient only includes expert outputs $E_{i}$ for the experts $i\in\mathcal{A}$ selected by the Top-K routing function. The other terms are essentially 0, which means that non-activated experts can not influence the router’s update for a token.
Our goal is to approximate the dense gradient in Equation˜4 without directly computing the missing terms $E_{i^{\prime}}(x)$ for $i^{\prime}\notin\mathcal{A}$. This enables the router to receive a signal from all experts in its gradient update, without actually activating all of these experts. The router can then consider information from all experts when learning to route tokens instead of being limited to only receiving feedback from experts that were activated. As a result, we expect an improved router which learns to allocate tokens to different experts more effectively, leading to improved MoE training performance.
Figure˜1 presents an overview of our method for updating the router. Since we wish to avoid additional computational overhead, we fill in the missing terms in Equation˜4 with estimates $\hat{E_{i}}$ for non-activated experts $E_{i}$. This simple addition ensures that the router updates its weights for all experts instead of only those selected by the Top-K routing.
@@ 3.1 Approximating Missing Expert Outputs
Backpropagating through the sparse MoE activation in Equation˜2 does not align with the router’s true dense gradient specified in Equation˜4. Specifically, the gradient in TopK routing leads to an error based on the non-activated experts. Let us consider the gradient of the router parameters $W$ with respect to the loss $\mathcal{L}$. For clarity, we examine the gradient term for a single input $x$; our approach also translates to practically training the router with a batch of inputs.
Using the chain rule, we decompose the router gradient to isolate the nondifferentiable $\frac{\partial y}{\partial\pi}$, similar to Equation˜3:
@@
The error term corresponding to $\frac{\partial y}{\partial\pi}$, $\sum_{i\notin\mathcal{A}}(E_{i}(x)-\mathbb{E}_{x}[E_{i}(x)])$, is 0 in expectation for all $i$. As a result, our default vector method corrects the gradient error by filling in missing expert activations with the mean expert output.
@@ 3.2 Default Expert Outputs with Exponential Moving Averages
To enable dense backpropagation without sacrificing the computational efficiency of sparse MoE forward passes, we propose Default MoE. Default MoE provides a default value for the outputs of non-selected experts using exponential moving averages (EMAs). The EMA accounts for the fact that experts are being updated, and thus their outputs for inputs $x$ will change over the course of training. For each expert $E_{i}$, we maintain an EMA of its average output:
@@
This approach enables dense backpropagation through the router while preserving the sparse computational pattern that makes MoE architectures efficient. The router receives gradient information about all possible routing decisions, not just the selected experts. We now evaluate DefaultMoE.
@@ 4 Evaluation
We describe the experimental setup in Section˜4.1, benchmark our Default MoE in Section˜4.2, show that our improvements hold across multiple configurations in Section˜4.3, discuss prior work in Section˜4.4, analyze our Default MoE in Section˜4.5, and find in Section˜4.6 that our method does not significantly impact throughput or the memory footprint of training, so the improvements we observe are -as far as we can tell- a free lunch. We conclude with a cohesive explanation of why Default MoE beats TopKMoE in Section˜5.
@@ 4.1 Experimental Setup
Model Architectures. We use NcK to refer to a model with $N$ total experts and $K$ active experts. We train both standard and finegrained (DeepSeek-AI Team, 2024a) MoEs. All of our models have 1.96 billion total parameters, 366 million of which are non-MoE parameters. This leaves 1.6B MoE parameters, and the number of active parameters depends on sparsity. All parameter counts are specified in Table˜1. We ablate the model architecture, hidden dim, number of total experts, and number of active experts in Section˜4.3.
@@
Dataset. We train on FineWeb-Edu (Lozhkov et al., 2024) (for Table˜2 and Figure˜12) and FineWeb (Penedo et al., 2024) (for all other plots and results) with the Llama3 tokenizer (Llama 3 Team, 2024). For our main results, we train for 160B tokens, which is a tokens-per-parameter ratio of $\approx 283$; at this level of overtraining, many spurious improvements should wash out, so we can be confident that any improvements we observe are real gains.
We have open-sourced our training code. All hyperparameters for the models trained in Table˜2 can be found in our config file, and all hyperparameters are the same between the baseline and our DefaultMoE. All further experimental details can be found in Section˜A.2.
@@ 4.2 Main Results
Benchmark 8c1 32c4
Random
Baseline
Score Improvement (%) Score Improvement (%)
TopK Default Diff (%) TopK Default Diff (%) TopK Default Diff (%) TopK Default Diff (%)
LogiQA 26.01.7 26.91.7 +3.6% 3.8 7.5 +96.0% 29.01.8 28.31.8 -2.6% 16.1 13.1 -19.0% 25.0
MathQA 26.20.8 25.80.8 -1.8% 4.9 3.0 -38.1% 25.00.8 26.10.8 +4.3% 0.1 4.4 +4266.7% 25.0
MMLU 31.80.4 32.30.4 +1.6% 27.3 29.4 +7.5% 33.70.4 32.50.4 -3.6% 34.7 29.9 -14.0% 25.0
OpenBookQA 37.22.2 38.22.2 +2.7% 48.8 52.8 +8.2% 39.62.2 39.02.2 -1.5% 58.4 56.0 -4.1% 25.0
Lambada 38.60.7 41.20.7 +6.6% 3761.8 4016.0 +6.8% 44.10.7 43.80.7 -0.6% 4307.1 4281.9 -0.6% 1.0111We use the ‘Random word from passage" as a baseline for LAMBADA (Paperno et al., 2016).
SocialIQA 39.71.1 41.01.1 +3.2% 20.3 24.2 +19.1% 41.11.1 40.71.1 -1.1% 24.7 23.3 -5.7% 33.0
HellaSwag 40.40.5 41.20.5 +2.0% 61.6 64.9 +5.4% 42.90.5 43.30.5 +0.9% 71.4 73.0 +2.2% 25.0
ARC 45.70.9 47.40.9 +3.7% 82.8 89.6 +8.3% 52.00.9 49.90.9 -4.2% 108.2 99.5 -8.1% 25.0
Winogrande 53.61.4 54.81.4 +2.2% 7.2 9.6 +33.0% 55.51.4 56.31.4 +1.4% 11.0 12.5 +14.4% 50.0
PubMedQA 55.22.2 52.42.2 -5.1% 67.3 58.8 -12.6% 54.42.2 57.42.2 +5.5% 64.8 73.9 +14.0% 33.0
BoolQ 58.50.9 62.00.8 +6.1% 17.0 24.1 +41.7% 62.80.8 63.10.8 +0.4% 25.6 26.2 +2.1% 50.0
PIQA 70.81.1 71.91.0 +1.5% 41.7 43.7 +5.0% 72.31.0 73.11.0 +1.1% 44.6 46.1 +3.4% 50.0
SciQ 86.21.1 87.81.0 +1.9% 244.8 251.2 +2.6% 89.11.0 90.30.9 +1.3% 256.4 261.2 +1.9% 25.0
Average 46.90.4 47.90.4 +2.1% 52.3 54.9 +5.0% 49.40.4 49.50.4 +0.3% 59.7 59.9 +0.4% -
Table 2: Comparison of Pretraining Benchmark Scores for TopK MoE and Default MoE. After training for 160 billion tokens, Default MoE outperforms TopK MoE across a range of benchmarks for both a standard MoE (right) and a finegrained MoE (left).
Pretraining Benchmarks. In Table˜2 we compare Default MoEs to TopK MoEs. All models are trained for 160B tokens on FineWeb-Edu and have 1.96B params. For both finegrained MoEs (32c4) and standard MoEs (8c1), our Default MoEs outperform TopK MoEs on standard benchmarks. We conduct all evaluations with the lm-eval harness (Gao et al., 2024).
Due to space constraints we defer pretraining curves to the Appendix, but discuss them here. In Figure˜12 we compare each MoE’s loss curves. Without introducing significant overhead, our method reduces the tokens required to reach a target perplexity of $12.18$ by $9\%$. As the model has just $500M$ active parameters, we are training for significantly more than the compute-optimal number of tokens Hoffmann et al. (2022). Therefore, we can be confident that we are not just seeing our method converge faster to a worse minima; this is a real improvement in terms of both speed of convergence and the quality of the converged model.
@@
In Table˜2 and Figure˜12 we pretrain on FineWeb-Edu (Lozhkov et al., 2024). We also pretrain models on FineWeb (Penedo et al., 2024), which is less curated. The rest of the plots in this paper (our ablations) are on FineWeb, in order to enforce that on some level we are not simply overfitting to the pretraining dataset that we use to collect benchmark scores. In Figure˜12 we show that when using the learning rate of $7\times 10^{-4}$ that we sweep in Figure˜2, our DefaultMoE outperforms the baseline TopK. Therefore, our improvements are not due to any unfair hyperparameter tuning when compared to the baseline.